Assume that a monopolist faces a demand curve for its product given by: p=100−1q p = 100 - 1 q Further assume that the firm's cost function is: TC=570+14q T C = 570 + 14 q Using calculus and formulas (don't just build a table in a spreadsheet as in the previous lesson) to find a solution, how much output should the firm produce at the optimal price? Round the optimal quantity to the nearest hundredth before computing the optimal price, which you should then round to the nearest cent. Note: Non-integer quantities may make sense when each unit of q represents a bundle of many individual items.
Answer #1

þ= 100-19 MR = 100-29 26 TC: 570+ 149 9 Mc-d(TC) 14 dq ing in the case of monopolistic condition firm is to produce at the le